Circular No. 6004

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993ak IN UGC 8685 C. Pollas, Observatoire de la Côte d’Azur, reports an independent

discovery a few months ago of this object (IAUC 6000) on a pair of films obtained on 1993 Mar. 15.08 UT at m_{v} ∼ 18.5; his private requests at that time to supernova search groups yielded no confirmation. He provides the following accurate position for SN 1993ak: α = 13^{h}40^{m}54^{s}.12, δ = +30^{o}35′16′′.9 (equinox 1950.0), which is 27′′.4 east and 4′′.2 south of the galaxy’s nucleus. A nearby stellar object of mag 18 has end figures 50^{s}.69, 34′54′′.5. The object is not visible on two films obtained 1994 Mar. 6.

COMET MUELLER (1993p) A. C. Gilmore, Mount John University Observatory, communicates

that a triple exposure of 5, 5, and 3 min on May 4.43 UT showed a diffuse, roughly parabolic patch—brighter at the apex—at the position of this object (cf. IAUC 5995). There was no hint of the small condensation that was present on Apr. 14.34 and well recorded in 1- and 2-min exposures. A double exposure of 10 and 15 min on May 5.38 and 5.39 confirmed the previous night’s images: a roughly parabolic envelope or fan ∼ 4′ long, with axis along p.a. 150^{o}, and no hint of condensation where a reasonably stellar condensation of mag 16–17 would have been recorded.

J. V. Scotti, Lunar and Planetary Laboratory, reports that observations on June 2.163 and 2.170 at large airmass by R. Jedicke with the Spacewatch Telescope revealed a diffuse and uncondensed image with no evidence of a nuclear condensation. The diffuse coma appeared as a roughly elliptical, nearly uniform image with its long axis aligned from p.a. 151^{o} to 331^{o} (dimensions ∼ 3′.5 × 1′.8). The total coma magnitude was near 10.5. The end of the ellipse toward p.a. 331^{o} appeared sharper than the opposing edge, giving the appearance of a comet with a broad, diffuse tail extending ∼ 4′ towards the south of the center of the elliptical image of the coma. Any remaining nucleus must have been fainter than V ∼ 18.5.

NOVA OPHIUCHI 1994 Corrigendum. A. Filippenko reports that further analysis of the spec-

trum of N Oph 1994 obtained by Van Dyk et al. (IAUC 6002) shows that the “forbidden lines” are actually permitted lines, mostly Fe II.

1994 June 8 Daniel W. E. Green